OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

the one that looks like \[\large y-\color{red}{y_1}=\color{blue}m(x-\color{green}{x_1})\] with \[\large \color{red}{y_1}=\color{red}{-1},\color{blue}m=\color{blue}4,\color{green}{x_1}=\color{green}{-3}\]

OpenStudy (anonymous):

thank you @satellite73 do you mind helping me with 2 more?

OpenStudy (anonymous):

not at all did you get the right answer for this one?

OpenStudy (anonymous):

yes @satellite73 here are the 2 last ones i need help with, they're kinda confusing. 1. Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20 2. Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

ok we start just as before, with \[y-(-3)=-6(x-(-2))\] or just \[y+3=-6(x+2)\] but this now you have to solve for \(x\) the steps are always the same distritute first on the right and get \[y+3=-6x-12\] then subtract \(3\) to get \[y=-6x-15\]

OpenStudy (anonymous):

can you do the second one?

OpenStudy (anonymous):

ok let me try one second @satellite73

OpenStudy (anonymous):

ok i got the first one but im having a bit trouble with the second one @satellite73